# Deflection angle in the strong deflection limit: a perspective from local geometrical invariants and matter distributions

### v2

@@ Abstract
where $\rho_{\mathrm{m}}$ and $\Pi_{\mathrm{m}}$ denote the local energy density and tangential pressure evaluated at the photon sphere of areal radius $R_{\mathrm{m}}$. This result reveals that $\bar{a}$ is intrinsically governed by the local matter distribution, with the universal value $\bar{a}=1$ emerging when $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}=0$. Notably, this finding resolves the long-standing puzzle of obtaining $\bar{a}=1$ in a class of spacetimes supported by a massless scalar field. Furthermore, these local properties are reflected in the frequencies of quasinormal modes, suggesting a profound connection between strong gravitational lensing and the dynamical response of gravitational wave signals. Our framework, independent of any specific gravitational theory, offers a universal tool for testing gravitational theories and interpreting astrophysical observations.
@@ II Photon dynamics and circular orbits in static, spherically symmetric spacetimes
We begin with the most general metric of a static, spherically symmetric spacetime, given by
@@
$m(r)=\frac{R}{2}\left[1-e^{-\lambda}(R^{\prime})^{2}\right].$ (4)
Furthermore, if $R^{\prime}\neq 0$, the metric function $e^{\lambda}$ can be expressed in terms of $m(r)$ as
$e^{\lambda(r)}=\frac{(R^{\prime})^{2}}{1-2m/R}\,,$ (5)
@@
Next, we examine photon dynamics in this spacetime. We assume that photons propagate along null geodesics; moreover, due to the spherical symmetry of the metric, their orbits can be confined to the equatorial plane (i.e., $\theta=\pi/2$) without loss of generality. Under these assumptions, the Lagrangian for a photon is given by
@@ III Derivation of the deflection angle in the strong deflection limit
which is different from the variable in the literature. This definition ensures that the final expressions are formulated primarily in terms of the areal radius. As a result, the formulation is both coordinate-invariant and geometrically meaningful. With this substitution, the integral $I(r_{0})$ becomes
@@ III.1 $R^{\prime}_{\mathrm{m}}\neq 0$
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{2}{\bar{a}^{2}}\left(1-\frac{2m(r_{\mathrm{m}})}{R_{\mathrm{m}}}\right)^{-1}\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi.$ (36)
@@
While previous works in the literature have expressed these coefficients arising from $I_{\mathrm{D}}$ in terms of metric functions, our formulation instead recasts them in terms of local quantities at the photon sphere (i.e., $V^{\prime\prime}_{\mathrm{m}}$ and $R_{\mathrm{m}}$) together with a well-defined quasi-local quantity (i.e., $m(r_{\mathrm{m}})$). This approach not only clarifies the physical origin of the logarithmic divergence in the deflection angle but also clearly identifies the observable quantities that govern the phenomenon in the strong deflection limit.
@@ III.2 $R^{\prime}_{\mathrm{m}}=0$
$\bar{b}$ $=\bar{a}\log\frac{4e^{\lambda_{\mathrm{m}}}}{\bar{a}^{2}R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi.$ (42)
@@
$\bar{b}$ $=\bar{a}\log\left[\>\!\frac{8}{\bar{a}^{2}}\left(1-2\>\!\frac{\mathrm{d}m}{\mathrm{d}R}\Big{|}_{r=r_{\mathrm{m}}}\right)^{-1}\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ (43)
where $\mathrm{d}m/\mathrm{d}R$ is an invariant measure of the local mass variation near the photon sphere.
@@ IV Formula for the deflection angle in the strong deflection limit
$\bar{b}=$ $\bar{a}\log\left[\frac{6}{\bar{a}^{2}\big{(}1+R_{\mathrm{m}}^{2}G_{(1)(1)}^{\mathrm{m}}\big{)}}\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ if $R^{\prime}_{\mathrm{m}}\neq 0$, (61a)
$\bar{b}=$ $\bar{a}\log\left[\frac{8}{\bar{a}^{2}\big{(}1-R_{\mathrm{m}}^{2}G_{(0)(0)}^{\mathrm{m}}\big{)}}\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ if $R^{\prime}_{\mathrm{m}}=0$. (61b)
@@
Consequently, these results imply that we can extract information about the curvature near the photon sphere by measuring the deflection angle in the strong deflection regime. Specifically, if $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed for direct comparison with observations—then we can deduce $G_{(1)(1)}^{\mathrm{m}}$ (with $m(r_{\mathrm{m}})$) determined via Eq. (55) in the case $R^{\prime}_{\mathrm{m}}\neq 0$, or $G_{(0)(0)}^{\mathrm{m}}$ (with $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$) determined via Eq. (56) in the case $R^{\prime}_{\mathrm{m}}=0$. In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
V Interpreting strong deflection in terms of local matter fields
In this section, we elucidate the relationship between the deflection angle in the strong deflection limit and matter field quantities. To this end, we adopt the following definitions:
@@ V Interpreting strong deflection in terms of local matter fields
where, in general relativity, $\rho(r)$ denotes the local energy density, $P(r)$ the radial pressure, and $\Pi(r)$ the tangential pressure as measured by static observers. In alternative gravitational theories, however, $\rho$, $P$, and $\Pi$ may be interpreted as components of a more general tensor encoding both matter distribution and additional curvature contributions. Note that, while we use the standard terminology for matter fields in general relativity, these expressions can be recast in terms of the corresponding curvature quantities via Eq. (62), thus extending the applicability of our discussion to other gravitational theories.
@@
$\bar{b}=$ $\bar{a}\log\left[\frac{6}{\bar{a}^{2}\left(1+8\pi R_{\mathrm{m}}^{2}P_{\mathrm{m}}\right)}\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ for $R^{\prime}_{\mathrm{m}}\neq 0$, (69a)
$\bar{b}=$ $\bar{a}\log\left[\frac{8}{\bar{a}^{2}\left(1-8\pi R_{\mathrm{m}}^{2}\rho_{\mathrm{m}}\right)}\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ for $R^{\prime}_{\mathrm{m}}=0$. (69b)
@@
In the following, we examine the relationship between the properties of the matter field quantities defined above and the strong field limit coefficients. Detailed calculations of the strong field limit coefficients for specific spacetime models have been extensively addressed in the literature, and we refer the reader to those works for further details. Nevertheless, it is instructive to present results for several well-known spacetimes. Furthermore, we elucidate the universal properties that emerge when particular combinations of the matter field quantities vanish.
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild metric
$\mathrm{d}s^{2}=-\left(1-\frac{2M}{r}\right)\mathrm{d}t^{2}+\left(1-\frac{2M}{r}\right)^{-1}\mathrm{d}r^{2}+r^{2}\>\!\mathrm{d}\Omega^{2}$ (70)
@@
Next, we consider an electrovacuum spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström metric
@@
$\bar{a}=\frac{1}{\sqrt{1-2Q^{2}/r_{\mathrm{m}}^{2}}},\quad\bar{b}=\bar{a}\log\left[\>\!\frac{6}{\bar{a}^{2}\left(1-Q^{2}/r_{\mathrm{m}}^{2}\right)}\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ (75)
@@
As a specific example, we consider a massless scalar field $\Phi$ minimally coupled to gravity. Its energy-momentum tensor is given by (see, e.g., Ref. [36])
@@
This result explicitly shows that $\rho+\Pi=0$ throughout the spacetime, which naturally leads to a universal behavior in the strong deflection limit by ensuring that $\bar{a}=1$. In other words, even though the spacetime is not strictly vacuum due to the presence of the massless scalar field, the specific structure of its energy-momentum tensor guarantees that $\bar{a}=1$, thereby retaining the Schwarzschild value.
This analysis clearly explains the property $\bar{a}=1$, which, although consistently observed in many disparate examples, had remained unexplained in terms of its fundamental physical origin. One notable example is the Janis–Newman–Winicour (JNW) spacetime [37, 38], which is described by the metric
$\mathrm{d}s^{2}=-\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{\gamma}\mathrm{d}t^{2}+\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{-\gamma}+\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{1-\gamma}r^{2}\>\!\mathrm{d}\Omega^{2},$ (78)
where $\gamma$ and $r_{\mathrm{g}}$ are constant parameters. For $11/2<\gamma<$ 1, this solution describes weakly naked singularities at $r=r_{\mathrm{g}}$, accompanied by a nontrivial scalar field profile, and features the photon sphere located at $r=r_{\mathrm{g}}(2\gamma+1)/2$. The associated scalar field quantities are given by
@@
$\bar{a}=1,\quad\bar{b}=\log\left[\>\!\frac{2(2\gamma+1)(2\gamma-1)}{\gamma^{2}}\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi$ (80)
@@
Another notable example is the Ellis–Bronnikov wormhole spacetime [39, 40], which is given by the metric
@@
$\alpha(b)$ $=-\bar{a}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)+\bar{b}+O\left(\left(\frac{b}{b_{\mathrm{c}}}-1\right)\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)\right)$ (83)
@@ VI Summary and discussion
$\omega_{\mathrm{QNM}}=\Omega_{\mathrm{c}}\>\!l-i\left(n+\frac{1}{2}\right)|\lambda_{\mathrm{L}}|$ (85)
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [42, 43], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [44].
Our formulation is a universal, coordinate-independent framework that applies to a broader range of spacetime models than conventional approaches, promising numerous future applications. Moreover, its independence from any specific gravitational theory enables its application to solutions beyond general relativity. Consequently, further investigations are required to clarify how the strong field limit coefficients depend on matter fields in modified gravitational theories, and our formulation is expected to serve as a foundation for clarifying their physical significance. Although the present study focuses on static, spherically symmetric spacetimes, extending the formulation to less symmetric cases remains an important direction for future research. In addition, applying our formulation to the strong deflection of massive particles offers another promising avenue for further investigation (for perturbations of the stable circular orbit of massive particles, see Ref. [41]). Furthermore, since a relationship between QNM frequencies and local curvature and matter fields is suggested, combining high-precision measurements of quasinormal modes from gravitational wave observations with analyses based on our formulation is expected to set the stage for a novel approach that yields detailed insights into the internal structure and local physical conditions of compact objects in extreme regimes.

### v3

@@ Abstract
where $\rho_{\mathrm{m}}$ and $\Pi_{\mathrm{m}}$ denote the local energy density and tangential pressure evaluated at the photon sphere of areal radius $R_{\mathrm{m}}$. This result reveals that $\bar{a}$ is intrinsically governed by the local matter distribution, with the universal value $\bar{a}=1$ emerging when $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}=0$. Notably, this finding resolves the long-standing puzzle of obtaining $\bar{a}=1$ in a class of spacetimes supported by a massless scalar field. Furthermore, these local properties are reflected in the frequencies of quasinormal modes, suggesting a profound connection between strong gravitational lensing and the dynamical response of gravitational wave signals.
@@ II Photon dynamics and circular orbits in static, spherically symmetric spacetimes
We begin with the most general line element of a static, spherically symmetric spacetime, given by
@@
$m(r)=\frac{R}{2}\left[1-e^{-\lambda}(R^{\prime})^{2}\right],$ (4)
where a prime denotes differentiation with respect to the radial coordinate $r$. Furthermore, if $R^{\prime}\neq 0$, the metric function $e^{\lambda}$ can be expressed in terms of $m(r)$ as
$e^{\lambda(r)}=\frac{(R^{\prime})^{2}}{1-2m(r)/R}\,$, (5)
@@
Next, we examine photon dynamics in this spacetime. We assume that photons propagate along null geodesics, that is, they interact only gravitationally with the background spacetime. This standard assumption is consistent with the conventional treatments of strong gravitational lensing in curved spacetimes (e.g., Refs. [5, 6]). In situations where light-matter interactions are non-negligible, such as in refractive or dispersive media (see, e.g., Refs. [30, 31]), the photon dynamics deviate from null geodesics, and the present formulation would need to be generalized accordingly.
Due to the spherical symmetry of the metric, their orbits can be confined to the equatorial plane (i.e., $\theta=\pi/2$) without loss of generality. Under these assumptions, the Lagrangian for a photon is given by
@@ III Derivation of the deflection angle in the strong deflection limit
which differs from the variable used in the literature. This definition ensures that the final expressions are formulated primarily in terms of the areal radius. As a result, the formulation is both coordinate-invariant and geometrically meaningful. With this substitution, the integral $I(r_{0})$ becomes
@@ III.1 $R^{\prime}_{\mathrm{m}}\neq 0$
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{2}\left(1-\frac{2m(r_{\mathrm{m}})}{R_{\mathrm{m}}}\right)\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi.$ (36)
@@
While previous works in the literature have expressed these coefficients arising from $I_{\mathrm{D}}$ in terms of metric functions, our formulation instead recasts them in terms of local quantities at the photon sphere (i.e., $V^{\prime\prime}_{\mathrm{m}}$ and $R_{\mathrm{m}}$) together with a well-defined quasi-local quantity (i.e., $m(r_{\mathrm{m}})$). This approach not only clarifies the physical origin of the logarithmic divergence in the deflection angle but also identifies the observable quantities governing the phenomenon in the strong deflection limit.
@@ III.2 $R^{\prime}_{\mathrm{m}}=0$
$\bar{b}$ $=-\bar{a}\log\frac{\bar{a}^{2}R_{\mathrm{m}}R^{\prime\prime}_{\mathrm{m}}}{4e^{\lambda_{\mathrm{m}}}}+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi.$ (42)
@@
$\bar{b}$ $=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{8}\left(1-2\>\!\frac{\mathrm{d}m}{\mathrm{d}R}\Big|_{r=r_{\mathrm{m}}}\right)\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi.$ (43)
Here $\mathrm{d}m/\mathrm{d}R$ denotes the variation of the quasi-local mass $m$ with respect to the areal radius $R$. Since $m$ is fundamentally defined as a function of the areal radius through the Misner–Sharp construction, $\mathrm{d}m/\mathrm{d}R$ provides an invariant measure of the local mass variation near the photon sphere.
@@ IV Formula for the deflection angle in the strong deflection limit
$\bar{b}=$ $-\bar{a}\log\left[\frac{\bar{a}^{2}}{6}\left(1+R_{\mathrm{m}}^{2}G_{(1)(1)}^{\mathrm{m}}\right)\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ if $R^{\prime}_{\mathrm{m}}\neq 0$, (61a)
$\bar{b}=$ $-\bar{a}\log\left[\frac{\bar{a}^{2}}{8}\left(1-R_{\mathrm{m}}^{2}G_{(0)(0)}^{\mathrm{m}}\right)\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ if $R^{\prime}_{\mathrm{m}}=0$. (61b)
@@
Consequently, these results imply that information about the curvature near the photon sphere can be extracted from measurements of the deflection angle in the strong deflection regime. Specifically, if $\bar{a}$ and $R_{\mathrm{m}}$ are determined from observations, the curvature combination $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$ can be directly derived. Moreover, if $\bar{b}$ is also measured and a specific spacetime model is assumed—so that all the relevant quantities can be computed for direct comparison with observations—then we can deduce $G_{(1)(1)}^{\mathrm{m}}$ (with $m(r_{\mathrm{m}})$) determined via Eq. (55) in the case $R^{\prime}_{\mathrm{m}}\neq 0$, or $G_{(0)(0)}^{\mathrm{m}}$ (with $\mathrm{d}m/\mathrm{d}R|_{r=r_{\mathrm{m}}}$) determined via Eq. (56) in the case $R^{\prime}_{\mathrm{m}}=0$. In other words, if the theoretical predictions for $\bar{a}$ and $\bar{b}$ agree reasonably well with the observed values, this approach enables us to infer the local curvature and the gravitational mass, thereby providing a robust consistency check of the underlying gravitational model.
V Interpreting strong deflection in terms of local matter fields in general relativity
In this section, we focus on general relativity, in which the Einstein equations directly relate the Einstein tensor components to the local matter variables—the energy density $\rho(r)$, the radial pressure $P(r)$, and the tangential pressure $\Pi(r)$—as
@@ V Interpreting strong deflection in terms of local matter fields in general relativity
where $\rho(r)$, $P(r)$, and $\Pi(r)$ denote the quantities measured by static observers. In this discussion, we do not assume that the matter sector is a perfect fluid, and hence no specific equation of state is imposed on the quantities $\rho(r)$, $P(r)$, and $\Pi(r)$. The term “matter” is used here in the standard general-relativistic sense to denote any contribution to the stress–energy tensor, including field configurations such as electromagnetic or scalar fields.
@@
$\bar{b}=$ $-\bar{a}\log\left[\frac{\bar{a}^{2}}{6}\left(1+8\pi R_{\mathrm{m}}^{2}P_{\mathrm{m}}\right)\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ for $R^{\prime}_{\mathrm{m}}\neq 0$, (69a)
$\bar{b}=$ $-\bar{a}\log\left[\frac{\bar{a}^{2}}{8}\left(1-8\pi R_{\mathrm{m}}^{2}\rho_{\mathrm{m}}\right)\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ for $R^{\prime}_{\mathrm{m}}=0$. (69b)
@@
In the following, we examine the relationship between the properties of the matter field quantities defined above and the strong field limit coefficients. Detailed calculations of the strong field limit coefficients for specific spacetime models have been extensively addressed in the literature, and we refer the reader to those works for further details. Nevertheless, it is instructive to present results for several well-known spacetimes. Furthermore, we elucidate the universal properties that emerge when specific combinations of matter-field quantities vanish.
As the simplest case, we consider a vacuum spacetime in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Schwarzschild line element
$\mathrm{d}s^{2}=-\left(1-\frac{2M}{r}\right)\mathrm{d}t^{2}+\left(1-\frac{2M}{r}\right)^{-1}\mathrm{d}r^{2}+r^{2}\>\!\mathrm{d}\Omega^{2},$ (70)
@@
Next, we consider an electrovacuum spacetime (i.e., a vacuum spacetime containing only electromagnetic fields) in the vicinity of the photon sphere. According to Birkhoff’s theorem, this region is described by the Reissner–Nordström line element
@@
$\bar{a}=\frac{1}{\sqrt{1-2Q^{2}/r_{\mathrm{m}}^{2}}},\quad\bar{b}=-\bar{a}\log\left[\>\!\frac{\bar{a}^{2}}{6}\left(1-Q^{2}/r_{\mathrm{m}}^{2}\right)\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ (75)
@@
As a specific example, we consider a massless scalar field $\Phi$ minimally coupled to gravity. Its stress-energy tensor is given by (see, e.g., Ref. [38])
@@
This result explicitly shows that $\rho+\Pi=0$ throughout the spacetime, which naturally leads to a universal behavior in the strong deflection limit by ensuring that $\bar{a}=1$. In other words, even though the spacetime is not strictly vacuum due to the presence of the massless scalar field, the specific structure of its stress-energy tensor guarantees that $\bar{a}=1$, thereby retaining the Schwarzschild value.
This analysis clearly explains the property $\bar{a}=1$, which, although consistently observed in many disparate examples, had remained unexplained in terms of its fundamental physical origin. One notable example is the Janis–Newman–Winicour (JNW) spacetime [39, 40], which is described by the line element
$\mathrm{d}s^{2}=-\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{\gamma}\mathrm{d}t^{2}+\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{-\gamma}\mathrm{d}r^{2}+\left(1-\frac{r_{\mathrm{g}}}{r}\right)^{1-\gamma}r^{2}\>\!\mathrm{d}\Omega^{2},$ (78)
where $\gamma$ and $r_{\mathrm{g}}$ are constant parameters. For $1/2<\gamma<1$, this solution describes weakly naked singularities at $r=r_{\mathrm{g}}$, accompanied by a nontrivial scalar field profile, and features the photon sphere located at $r=r_{\mathrm{g}}(2\gamma+1)/2$. The associated scalar field quantities are given by
@@
$\bar{a}=1,\quad\bar{b}=\log\left[\>\!\frac{2(2\gamma+1)(2\gamma-1)}{\gamma^{2}}\>\!\right]+I_{\mathrm{R}}(r_{\mathrm{m}})-\pi,$ (80)
@@
Another notable example is the Ellis–Bronnikov wormhole spacetime [41, 42], which is given by the line element
@@
$\alpha(b)$ $=-\bar{a}\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)+\bar{b}+O\left(\left(\frac{b}{b_{\mathrm{c}}}-1\right)\log\left(\frac{b}{b_{\mathrm{c}}}-1\right)\right),$ (83)
@@ VI Summary and discussion
$\omega_{\mathrm{QNM}}=\Omega_{\mathrm{c}}\>\!l-i\left(n+\frac{1}{2}\right)|\lambda_{\mathrm{L}}|,$ (85)
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. [^4] It should be noted that the correspondence between photon spheres and eikonal QNMs is not strictly universal; for example, in Lovelock gravity, the eikonal QNMs frequencies of gravitational perturbations deviate from the photon-sphere prediction due to the spin-dependent structure of the effective potential [44]. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [45, 46], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [47].
Our formulation is a universal, coordinate-independent framework that applies to a broader range of spacetime models than conventional approaches, promising numerous future applications. Moreover, when expressed in terms of the Einstein tensor, the formulation becomes independent of any specific gravitational theory, allowing its application to solutions beyond general relativity. Consequently, further investigations are required to clarify how the strong field limit coefficients depend on matter fields in modified gravitational theories, and our formulation is expected to serve as a foundation for clarifying their physical significance. Although the present study focuses on static, spherically symmetric spacetimes, extending the formulation to less symmetric cases remains an important direction for future research. In addition, applying our formulation to the strong deflection of massive particles offers another promising avenue for further investigation (for perturbations of the stable circular orbit of massive particles, see Ref. [43]). Furthermore, since a relationship between QNM frequencies and local curvature and matter fields is suggested, combining high-precision measurements of quasinormal modes from gravitational wave observations with analyses based on our formulation is expected to set the stage for a novel approach that yields detailed insights into the internal structure and local physical conditions of compact objects in extreme regimes.